# Uniform convergence of the smooth calibration error and its relationship with functional gradient

### v2

@@ Appendix E Proofs for the kernel boosting
Since $\phi\in\mathcal{H}$ implies $-\phi\in\mathcal{H}$, we define $\tilde{\phi}:=-\phi$. Then, the margin assumption implies:
@@ E.2 Proof of Theorem 4
$\sum_{t=0}^{T-1}\frac{1}{2}w_{t}\|\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}\leq L(g^{(0)}).$ (206)
In particular, using constant step size $w_{t}=\eta$, we obtain:
$\frac{1}{T}\sum_{t=0}^{T-1}\|\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}\leq\frac{2}{\eta T}L(g^{(0)}).$ (207)
From Eq. (194), we have:
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla_{g}\ell(g^{(t)}(X),Y)\right\|_{L_{1}(S_{n})}^{2}\leq\frac{2}{\eta T}L(g^{(0)}).$ (208)
@@
$\left\|\nabla_{g}\ell(\bar{g}^{(T)}(X),Y)\right\|_{L_{1}(S_{n})}\leq\sqrt{\frac{2}{\eta T}L(g^{(0)})}.$ (210)
@@
We now evaluate the Rademacher complexity. Since $\mathcal{F}=\sigma\circ\mathcal{H}$ and $\sigma$ is $1/4$-Lipschitz, Talagrand’s contraction lemma yields:
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})=\mathfrak{R}_{\mathcal{D},n}(\sigma\circ\mathcal{H})\leq\frac{1}{4}\mathfrak{R}_{\mathcal{D},n}(\mathcal{H})\leq\frac{\alpha}{4}\sqrt{\frac{\Lambda}{n}},$ (211)
where $\alpha$ is an upper bound on the RKHS norm $\|g\|_{\mathcal{H}}\leq\alpha$, which we estimate below.
@@
$\|g^{(T)}\|_{\mathcal{H}}$ $\leq\|g^{(0)}\|_{\mathcal{H}}+\left\|\eta\sum_{t=0}^{T-1}\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\right\|_{\mathcal{H}}.$ (213)
@@
$\|\eta\sum_{t=0}^{T-1}\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}\leq\frac{1}{T}\sum_{t=0}^{T-1}\|\eta T\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}$ $\leq\eta^{2}T^{2}\frac{1}{T}\sum_{t=0}^{T-1}\|\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}$ (214)
$\leq\eta^{2}T^{2}\frac{2}{\eta T}L(g^{(0)})$ (215)
$\leq 2\eta TL(g^{(0)})$ (216)
@@
$\|g^{(T)}\|_{\mathcal{H}}\leq\Lambda^{\prime}+\sqrt{2\eta TL(g^{(0)})}.$ (217)
Absorbing $\sqrt{2L(g^{(0)})}$ into a universal constant, we conclude:
$\alpha=\|g^{(T)}\|_{\mathcal{H}}=\mathcal{O}(\Lambda^{\prime}+\sqrt{\eta T}).$
Substituting this into the Rademacher bound yields the final generalization result.
E.3 Proof of Misclassification rate
@@
$P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq\frac{(1+e^{\rho})}{\gamma}\sqrt{\frac{L_{n}(g^{(0)})}{\eta T}}+\frac{2(\Lambda^{\prime}+\sqrt{2\eta TL(g^{(0)})})}{\rho}\sqrt{\frac{\Lambda}{n}}+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}.$ (220)
This result illustrates a trade-off with respect to $\eta T$, which balances the training misclassification rate and the complexity of the hypothesis class.
E.4 Proof of Corollary 4
We first remark that both the smooth calibration error and the misclassification rate exhibit similar asymptotic behavior with respect to $n$ and $\eta T$. By substituting the given hyperparameters into the bound presented in Theorem 4, we obtain the desired result.
Appendix F Formal settings of the two-layer neural network
F.1 Formal results for the smooth CE
@@
The result above corresponds to Eq. (26). This bound provides a guarantee on the training smooth CE. Proposition 8 in nitanda2019gradient yields estimates for the Rademacher complexity and covering numbers.
@@
^$\hat{\mathfrak{R}}_{S}(\mathcal{F})$ $\leq Cm^{1/2-\beta}D_{w,m,T}(1+K_{1}+K_{2})$
@@
^$\hat{\mathfrak{R}}_{S}(\mathcal{F})$ $\leq\frac{8K_{1}m^{1/2-\beta}}{\sqrt{n}}\left(D_{w,m,T}+\sqrt{\frac{\log(Am/\delta)}{b}}\right).$ (224)
@@
$\qquad\cdot\sqrt{\frac{d}{n}\log\left(n(1+K_{1}+K_{2})\left(\log(m/\delta)+D_{w,m,T}^{2}\right)\right)})+\frac{3+\sqrt{2}}{2}\sqrt{\frac{\log\frac{5}{\delta}}{n}}.$ (226)
@@
$\frac{C_{2}}{\sqrt{n}}+\frac{C_{4}K_{1}m^{1/2-\beta}}{\sqrt{n}}\left(D_{w,m,T}+\sqrt{\frac{\log(Am/\delta)}{b}}\right)+\frac{3+\sqrt{2}}{2}\sqrt{\frac{\log\frac{5}{\delta}}{n}}.$ (228)
@@
We use the following upper bound
$\mathfrak{R}_{\mathcal{D},2n}(\mathcal{F})\leq\hat{\mathfrak{R}}_{S}(\mathcal{F})+\sqrt{\frac{\log\frac{2}{\delta}}{4n}}$ (229)
where $S$ is the size $2n$ dataset drawn i.i.d from $\nu$. Then we set $\delta\to\frac{2}{5}\delta$ in Eq. (229), and $\delta\to\frac{3}{5}\delta$ in Theorem 2. Then, the sum of the confidence terms becomes:
$\sqrt{\frac{\log\frac{2}{\frac{2}{5}\delta}}{4n}}+\left(1+\frac{1}{\sqrt{2}}\right)\sqrt{\frac{\log\frac{3}{\frac{3}{5}\delta}}{n}}=\left(1+\frac{1}{2}\frac{1}{\sqrt{2}}\right)\sqrt{\frac{\log\frac{5}{\delta}}{n}}=\frac{3+\sqrt{2}}{2}\sqrt{\frac{\log\frac{5}{\delta}}{n}}$ (230)
For completeness, we include the result on the misclassification error, adapted from Theorem 4 in nitanda18a:
@@
F.2 Formal statement for Corollary 5
@@
Appendix G Relationships of the proper scoring rule and functional gradient
@@
From Eqs. (235) and (239), we obtain upper bounds on the training smooth CE in terms of the functional gradient evaluated on the training dataset:
@@ Appendix H Experiment
H.1 Gradient Boosting Trees
@@
$X\mid Y=0\sim N\left(\begin{bmatrix}-1.3\\ -1\end{bmatrix},\begin{bmatrix}(1.2\cdot 1.3)^{2}&0\\ 0&1.2^{2}\end{bmatrix}\right),$ (245)
$X\mid Y=1\sim N\left(\begin{bmatrix}1\\ 1.3\end{bmatrix},\begin{bmatrix}1.2^{2}&0\\ 0&(1.2\cdot 1.3)^{2}\end{bmatrix}\right).$ (246)
@@
Figure 1: GBT experiments on the toy dataset defined in Eq. (246) with $m\geq d$
@@
Figure 2: GBT experiments on the toy dataset defined in Eq. (246) with $m<d$
@@
H.2 Numerical experiments over two-layer neural network
We conducted similar experiments using a two-layer neural network corresponding to Section 4.3. We prepared the toy dataset and designed the two-layer neural network so that Assumption 1 in Appendix F is satisfied. We used the sigmoid activation function, initialized the parameters independently from the standard Gaussian distribution, and set the number of hidden units to be even to satisfy Assumption (A3). The coefficients $a_{r}$ were also set according to the specification in Assumption (A3). Unless otherwise stated, we used 300 hidden units. For optimization, we used full-batch gradient descent (GD) with a fixed step size of $w=0.01$.
First, we performed experiments on the same toy dataset used for GBTs and evaluated the training and test metrics by increasing the number of GD iterations while fixing the training dataset size ($n=200$). Figure 4(a) reports the same set of metrics as in the GBT experiments (Figure 1). We observed that the smooth CE decreases monotonically, consistent with the upper bound behavior predicted by Eq. (26). However, from the middle and right panels of Figure 4(a), we observe that the test smooth CE does not decrease as the number of training iterations increases. We also confirmed that the generalization gap of the cross-entropy loss increases as $T$ increases. According to Eq. (225), the complexity term grows with the number of iterations, which may explain this phenomenon. On the other hand, we did not observe a clear increase in the generalization gap of the smooth CE as $T$ increased.
Next, we varied the training dataset size while setting the number of iterations to $T=\mathcal{O}(\sqrt{n})$, and evaluated various metrics. The results are shown in Figure 4(b). From the middle and right panels, we observe that both the test smooth CE and the generalization gap decrease monotonically as $n$ increases. This is consistent with the theoretical discussion provided in Appendix F.
@@ H.2 Numerical experiments over two-layer neural network
Figure 4: Two-layer neural network using toydata defined in Eq. (246)
@@
Figure 5: Two-layer neural network using toy dataset defined in Eq. (247)

### v4

@@ Appendix F Proofs for the kernel boosting
Since $\phi\in\mathcal{H}$ implies $\phi\in\mathcal{H}$, we define $\tilde{\phi}:=-\phi$. Then, the margin assumption implies:
@@ F.2 Proof of Theorem 4
$\sum_{t=0}^{T-1}\frac{1}{2}w_{t}\|\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}\leq L_{n}(g^{(0)}).$ (238)
In particular, using constant step size $w_{t}=w$, we obtain:
$\frac{1}{T}\sum_{t=0}^{T-1}\|\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}\leq\frac{2}{wT}L_{n}(g^{(0)}).$ (239)
From Eq. equation 226, we have:
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla_{g}\ell(g^{(t)}(X),Y)\right\|_{L_{1}(S_{n})}^{2}\leq\frac{2}{wT}L_{n}(g^{(0)}).$ (240)
@@
$\left\|\nabla_{g}\ell(\bar{g}^{(T)}(X),Y)\right\|_{L_{1}(S_{n})}\leq\sqrt{\frac{2}{wT}L_{n}(g^{(0)})}.$ (242)
@@
We now evaluate the Rademacher complexity. We express the function class of $\bar{g}^{(T)}=\frac{1}{n}\sum_{t=0}^{T-1}g^{t}$ as $\mathcal{G}$. Since $\mathcal{F}=\sigma\circ\mathcal{G}$ and $\sigma$ is $1/4$-Lipschitz, Talagrand’s contraction lemma yields:
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})=\mathfrak{R}_{\mathcal{D},n}(\sigma\circ\mathcal{G})\leq\frac{1}{4}\mathfrak{R}_{\mathcal{D},n}(\mathcal{G})\leq\frac{\alpha}{4}\sqrt{\frac{\Lambda}{n}},$ (243)
where $\alpha$ is an upper bound on the RKHS norm of $\mathcal{G}$, which we estimate below. The final inequality follows from the standard Rademacher complexity estimate of the RKHS, see Mohri for the details.
@@
$\|g^{(T)}\|_{\mathcal{H}}$ $\leq\|g^{(0)}\|_{\mathcal{H}}+\left\|w\sum_{t=0}^{T-1}\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\right\|_{\mathcal{H}}.$ (245)
@@
$\|w\sum_{t=0}^{T-1}\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}\leq\frac{1}{T}\sum_{t=0}^{T-1}\|wT\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}$ $\leq w^{2}T^{2}\frac{1}{T}\sum_{t=0}^{T-1}\|\mathcal{T}_{k}\nabla_{g}L_{n}(g^{(t)})\|_{\mathcal{H}}^{2}$ (246)
$\leq w^{2}T^{2}\frac{2}{wT}L_{n}(g^{(0)})$ (247)
$\leq 2wTL_{n}(g^{(0)})$ (248)
@@
$\|g^{(T)}\|_{\mathcal{H}}\leq\Lambda^{\prime}+\sqrt{2wTL_{n}(g^{(0)})}.$ (249)
Absorbing $\sqrt{2L_{n}(g^{(0)})}$ into a universal constant. When considering the norm of $\bar{g}^{(T)}$, it is also $\mathcal{O}(\Lambda^{\prime}+\sqrt{wT})$. Thus, we conclude $\alpha=\mathcal{O}(\Lambda^{\prime}+\sqrt{wT})$. Substituting this into the Rademacher bound yields the final generalization result.
F.3 Proof of Misclassification rate
@@
$P_{(X,Y)\sim\mathcal{D}}[(2Y-1)\bar{g}^{(T)}(X)\leq 0]\leq\frac{(1+e^{\rho})}{\gamma}\sqrt{\frac{L_{n}(g^{(0)})}{wT}}+\frac{2(\Lambda^{\prime}+\sqrt{2wTL_{n}(g^{(0)})})}{\rho}\sqrt{\frac{\Lambda}{n}}+\sqrt{\frac{\log\frac{1}{\delta}}{2n}}.$ (252)
This result illustrates a trade-off with respect to $wT$, which balances the training misclassification rate and the complexity of the hypothesis class.
F.4 Proof of Corollary 4
We first remark that both the smooth calibration error and the misclassification rate exhibit similar asymptotic behavior with respect to $n$ and $wT$. By substituting the given hyperparameters into the bound presented in Theorem 4, we obtain the desired result.
Appendix G Formal settings of the two-layer neural network
G.1 Formal results for the smooth CE
@@
The result above corresponds to Eq. equation 27. This bound provides a guarantee on the training smooth CE. Proposition 8 in nitanda2019gradient yields estimates for the Rademacher complexity and covering numbers.
@@
$\hat{\mathfrak{R}}_{S}(\mathcal{F})$ $\leq Cm^{1/2-\beta}D_{w,m,T}(1+K_{1}+K_{2})$
@@
$\hat{\mathfrak{R}}_{S}(\mathcal{F})$ $\leq\frac{8K_{1}m^{1/2-\beta}}{\sqrt{n}}\left(D_{w,m,T}+\sqrt{\frac{\log(Am/\delta)}{b}}\right).$ (256)
@@
$\qquad\cdot\sqrt{\frac{d}{n}\log\left(n(1+K_{1}+K_{2})\left(\log(2m/\delta)+D_{w,m,T}^{2}\right)\right)})+6\sqrt{\frac{\log\frac{6}{\delta}}{n}}.$ (258)
@@
$\frac{C_{2}}{\sqrt{n}}+\frac{C_{4}K_{1}m^{1/2-\beta}}{\sqrt{n}}\left(D_{w,m,T}+\sqrt{\frac{\log(2Am/\delta)}{b}}\right)+6\sqrt{\frac{\log\frac{6}{\delta}}{n}}.$ (260)
@@
We use the following upper bound from Mohri: with probability $1-\delta/2$, we have
$\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})\leq\hat{\mathfrak{R}}_{S}(\mathcal{F})+\sqrt{\frac{\log\frac{2}{\delta}}{2n}}.$ (261)
Then from Theorem 2 with probability $1-\delta$, we have
$\sup_{f\in\mathcal{F}}|\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$ $\leq\frac{C_{2}}{\sqrt{n}}+4\hat{\mathfrak{R}}_{S}(\mathcal{F})+6\sqrt{\frac{\log\frac{3}{\delta}}{n}}.$ (262)
We then substitute Eq. equation 7 and equation 256 and taking the union bound, we obtain the result. ∎
For completeness, we include the result on the classification error, adapted from Theorem 4 in nitanda18a:
@@
G.2 Formal statement for Corollary 5
@@
Appendix H Relationships of the proper scoring rule and functional gradient
@@
From Eqs. equation 267 and equation 271, we obtain upper bounds on the training smooth CE in terms of the functional gradient evaluated on the training dataset:
@@ Appendix I Experiment
I.1 Gradient Boosting Trees
@@
[…]
@@
Figure 1: GBT experiments on the toy dataset defined in Eq. equation 278 with $m\geq d$
@@
Figure 2: GBT experiments on the toy dataset defined in Eq. equation 278 with $m<d$
@@
I.2 Numerical experiments over two-layer neural network
We conducted similar experiments using a two-layer neural network corresponding to Section 4.3. We prepared the toy dataset and designed the two-layer neural network so that Assumption 1 in Appendix G is satisfied. We used the sigmoid activation function, initialized the parameters independently from the standard Gaussian distribution, and set the number of hidden units to be even to satisfy Assumption (A3). The coefficients $a_{r}$ were also set according to the specification in Assumption (A3). Unless otherwise stated, we used 300 hidden units. For optimization, we used full-batch gradient descent (GD) with a fixed step size of $w=0.01$.
First, we performed experiments on the same toy dataset used for GBTs and evaluated the training and test metrics by increasing the number of GD iterations while fixing the training dataset size ($n=200$). Figure 4(a) reports the same set of metrics as in the GBT experiments (Figure 1). We observed that the smooth CE decreases monotonically, consistent with the upper bound behavior predicted by Eq. equation 27. However, from the middle and right panels of Figure 4(a), we observe that the test smooth CE does not decrease as the number of training iterations increases. We also confirmed that the generalization gap of the cross-entropy loss increases as $T$ increases. According to Eq. equation 257, the complexity term grows with the number of iterations, which may explain this phenomenon. On the other hand, we did not observe a clear increase in the generalization gap of the smooth CE as $T$ increased.
Next, we varied the training dataset size while setting the number of iterations to $T=\mathcal{O}(\sqrt{n})$, and evaluated various metrics. The results are shown in Figure 4(b). From the middle and right panels, we observe that both the test smooth CE and the generalization gap decrease monotonically as $n$ increases. This is consistent with the theoretical discussion provided in Appendix G.
@@ I.2 Numerical experiments over two-layer neural network
Figure 4: Two-layer neural network using toydata defined in Eq. equation 278
@@
Figure 5: Two-layer neural network using toy dataset defined in Eq. equation 279